OpenStudy (anonymous):

Find all numbers x such that \[x + 3^x \lt 4\]Find all numbers x such that \[x + 3^x \lt 4\]@Mathematics

OpenStudy (anonymous):

I'm stuck on this one and the next few... :(

OpenStudy (zarkon):

notice that when x=1 \[x+3^x=4\]

OpenStudy (anonymous):

yeah, but what if I don't notice?

OpenStudy (anonymous):

like in the next few questions?

OpenStudy (zarkon):

if you had something other than 4 you would probably have to solve numerically.

OpenStudy (anonymous):

graph them. it's true for x < 1

OpenStudy (anonymous):

graph y = x + 3^x and y = 4

OpenStudy (anonymous):

no graphing calculators existed in 1969

OpenStudy (anonymous):

you don't need a calculator to draw those two graphs

OpenStudy (anonymous):

well..... I don't want to use graphs because we are still using just numbers

OpenStudy (anonymous):

i see no way of solving 3^x + x = 4 by elementary methods. simply guessing x = 1 is the only way. you could possibly use Lambert W function

OpenStudy (anonymous):

I guess I'll have to use what works then :-(

OpenStudy (anonymous):

sorry agdgdgdgwngo i have one more problem for you to answer plz help me???

OpenStudy (anonymous):

x = 4 + W/ln(3), where W = Lambert(3^-4/ln(3)) = 1

OpenStudy (anonymous):

x = 4 + W/ln(3), where W = Lambert(1/(27ln(3))) = 1

OpenStudy (anonymous):

oops: x = 4 + W/ln(3), where W = Lambert(1/(81ln(3))) = 1

OpenStudy (anonymous):

oops got it wrong again: 3^x + x = 4 3^x = 4 - x (4 - x)*3^(-x) = 1 Let m = 4 - x, then x = 4 - m m * 3^(m-4) = 1 thus m * 3^(m) * 3^(-4) = 1 thus m * 3^m = 81 so m * e^(ln(3)*m) = 81 so ln(3)*m * e^(ln(3)*m) = 81*ln(3) hence ln(3)*m = W(81*ln(3)) thus m = W(81*ln(3))/ln(3) thus x = 4 - W(81*ln(3))/ln(3), where W(x) is Lambert W function = 1